# SLAM3R: Real-Time Dense Scene Reconstruction from Monocular RGB Videos

### v2

@@ Appendix B Details for experimental settings
On the Replica dataset, we reconstruct the entire scene geometry using all video frames. The scene is initialized with a 5-frame window, and the local pointmap is recovered for each frame using the I2P model with a window size of 11. Within each window, frames are sampled at intervals of $I=20$ to ensure reasonable camera motion. By default, we co-register $K=10$ keyframes at a time, which share the same $S=10$ scene frames as a reference. The $S$ scene frames are selected through a two-step process. First, we calculate the correlation score between all frames in the buffering set and the $K$ keyframes. Then, we select $S$ frames from the buffering set that show the highest total correlation score with these keyframes. After every $R=20$ registered keyframes, we update the buffering set by retaining the keyframes with the highest reconstruction scores, where reconstruction score of a frame is the product of its mean confidence predicted by I2P and L2W model.
@@
Following recent research [59], we uniformly sample frames from the 7 Scenes dataset before evaluation. The frames are sampled with a stride of 20 from each test sequence, and we only reconstruct the points from the sampled frames. To handle sampled-frame-only input, we adapt our reconstruction pipeline for full-video input by setting $I=1$, $K=2$, $S=5$, and $R=1$ in practice.
@@
The global optimization with complete graph setting in DUSt3R and MASt3R requires substantial GPU memory. Consequently, to evaluate the global reconstruction quality of these two methods on the Replica dataset, we uniformly sample 1/20 of the images. As MASt3R provides model weights trained exclusively on a resolution of 512px, we specifically use images with resolutions of 512$\times$384 and 512$\times$288 as inputs for reconstructing the 7 Scenes [49] and Replica [52] datasets, respectively. In contrast, our method center-crops and resizes the input images to a fixed resolution of 224$\times$224, which results in less overlap between adjacent frames, making reconstruction inherently more challenging.
@@
Appendix C Additional analyses
@@
Camera tracking results.
Our method is designed in a new paradigm that reconstructs 3D points end-to-end without explicitly solving camera parameters. We have found that camera poses and scene reconstruction results are not fully positively correlated. Here, following NICER-SLAM [77], we evaluate the camera pose accuracy on 7 Scenes [49] and Replica [52] datasets using absolute trajectory error (ATE-RMSE). It is calculated by aligning the predicted camera trajectory with the GT trajectory. The predicted trajectory is derived by estimating camera poses through the PnP-RANSAC solver in OpenCV [6], leveraging the predicted pointmaps and GT intrinsics of each frame. Camera pose translations are then used to construct the trajectory. For DUSt3R [62] and MASt3R [26], we evaluate both the camera poses generated during global optimization and those derived via the PnP-RANSAC solver with their predicted pointmaps and GT intrinsic parameters. When evaluating Spann3R [59] on the Replica [52] dataset, only one-twentieth of the frames are used, as it fails to give reasonable results with all frames input.
The results are presented in Table 5 and Table 6. We outperform the concurrent work Spann3R [59], demonstrating the effectiveness of our hierarchical design with multi-view input and global retrieval. Among classical SLAM systems, the pose errors of GO-SLAM [73] and DROID-SLAM [54] are lower than those of NICER-SLAM. However, their reconstruction accuracy and completeness are worse. A similar pattern appears in comparison with DUSt3R — while its poses are worse than NICER-SLAM, it achieves the best reconstruction quality on the Replica dataset. Therefore, pose accuracy and reconstruction quality sometimes show inconsistencies. This discrepancy between pose and reconstruction errors indicates that effective end-to-end 3D reconstruction is possible and promising without first obtaining precise camera poses.

### v3

@@ Appendix B Details for experimental settings
Calculation of the evaluation metrics.
To evaluate reconstruction quality, we use accuracy and completeness as our metrics. They are calculated by:
$Accuracy=\frac{1}{P}\sum_{i=1}^{P}min_{j}(D(x_{i},y_{j})),$
$Completeness=\frac{1}{Q}\sum_{j=1}^{Q}min_{i}(D(x_{i},y_{j})).$
$P$ and $Q$ are the numbers of points in the reconstructed point cloud and GT point cloud respectively. $D(\cdot)$ represents Euclidean distance, and $x_{i}$ and $y_{j}$ represent iterating each point from the reconstructed and GT point cloud.
To measure the efficiency, we report FPS (frames per second), which is calculated by:
$FPS=F/time,$
where $time$ is the total time used to reconstruct the scene, and $F$ is the number of frames from the video.
We evaluate the camera pose accuracy using absolute trajectory error (ATE-RMSE), which is formulated by:
$ATE\mbox{-}RMSE=\sqrt{\frac{1}{F}\sum_{i=1}^{F}D(T^{gt}_{i},T^{perd}_{i})^{2}},$
where $T^{perd}$ and $T^{gt}$ are the camera center positions of the predicted and GT camera trajectories.
@@
On the Replica dataset, we reconstruct the entire scene geometry using all video frames. With the stride of the sliding window set to 1, all frames will be used as a keyframe once. For each window, frames are sampled around the keyframe, with $Skip=20$ frames per supporting frame, to ensure reasonable camera motion (disparity). We co-register $Co=10$ keyframes at each time, which share the same $K=10$ scene frames as a reference. These scene frames are selected through a two-step process. First, we calculate the correlation score between all frames in the buffering set and the $Co$ keyframes. Then, we select $K$ frames from the buffering set that show the highest total correlation score with these keyframes. After every $R=20$ registered keyframes, we update the buffering set by retaining the keyframes with the highest reconstruction scores, where reconstruction score of a frame is the product of its mean confidence predicted by I2P and L2W model. The insertion/update follows the reservoir sampling probability described in the main paper.
@@
Following Spann3R [61], the frames in each test sequence are sampled with a stride of 20, and we only reconstruct the points from the sampled frames. To handle sampled-frame-only input, we adapt our reconstruction pipeline for full-video input by setting $Skip=1$, $Co=2$, $K=5$, and $R=1$ in practice.
@@
The global optimization with complete graph setting in DUSt3R and MASt3R requires substantial GPU memory. Consequently, to evaluate the global reconstruction quality of these two methods on the Replica dataset, we uniformly sample 1/20 of the images. DUSt3R is tested using the weight-224 model with a resolution of 224$\times$224, the same as our input resolution, while MASt3R is tested using the weight-512 model with resolutions of 512$\times$384 and 512$\times$288 as inputs for reconstructing the 7 Scenes [51] and Replica [54] datasets, respectively. Note that a resolution of 224$\times$224 results in less overlap between adjacent frames, making reconstruction inherently more challenging.
@@
Appendix C Additional comparisons and analyses
More numerical results.
We report more quantitative comparisons of reconstruction results on ScanNet [11], Tanks and Temples [26], and ETH3D [50] datasets. We sampled three scenes from each dataset, and report the results in Table 8. SLAM3R outperforms Spann3R in most cases and demonstrates performance either comparable to or better than DUSt3R. These results further verify our method’s effectiveness.
ScanNet scene0011_00 scene0015_00 scene0019_00 Average
DUSt3R [64] 5.56 / 3.76 5.04 / 4.10 4.52 / 4.74 5.04 / 4.20
Spann3R [61] 13.09 / 11.37 8.51 / 7.79 7.97 / 9.66 9.86 / 9.61
SLAM3R (Ours) 5.86 / 3.98 5.98 / 5.97 4.27 / 4.34 5.37 / 4.76
Tanks and Temples Ignitius Truck Caterpillar Average
DUSt3R [64] 3.55/ 1.22 9.31 / 4.85 12.67 / 5.25 8.51 / 3.77
Spann3R [61] 5.51 / 1.10 6.40 / 12.61 11.50/ 5.74 7.80 / 6.48
SLAM3R (Ours) 3.30 / 0.94 5.35/ 5.59 12.26 / 5.05 6.97 / 3.86
ETH3D plant_scene_1 table_3 sofa_1 Average
DUSt3R [64] 2.98 / 2.48 3.13 / 1.30 2.05 / 3.67 2.72 / 2.48
Spann3R [61] 2.54 / 4.25 3.03 / 2.08 2.10 / 4.55 2.56 / 3.62
SLAM3R (Ours) 2.36 / 1.98 2.75 / 1.34 2.13 / 1.90 2.41 / 1.74
Table 8: Reconstruction errors (accuracy / completeness) on ScanNet [11], Tanks and Temples [26], and ETH3D [50] datasets.
@@ Appendix C Additional comparisons and analyses
Camera pose estimation.
The detailed results are presented in Table 6 and Table 7. For DUSt3R [64] and MASt3R [28], we evaluate the camera poses derived via the PnP-RANSAC solver with their predicted pointmaps (after global alignment) and GT intrinsic parameters. When evaluating Spann3R [61] on the Replica [54] dataset, only one-twentieth of the frames are used, as it fails to give reasonable results with all frames input.
We outperform the concurrent work Spann3R [61], demonstrating the effectiveness of our hierarchical design with multi-view input and global retrieval. Among classical SLAM systems, the pose errors of GO-SLAM [75] and DROID-SLAM [56] are lower than those of NICER-SLAM. However, their reconstruction accuracy and completeness are worse. This discrepancy between pose and reconstruction errors indicates that effective end-to-end 3D reconstruction is possible and promising without first obtaining precise camera poses.